OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 10x = 15

OpenStudy (anonymous):

25..

Parth (parthkohli):

\( \color{Black}{\Rightarrow ( a+ b)^2 = a^2 + 2ab + b^2}\) right?

OpenStudy (anonymous):

aree you sure??

OpenStudy (anonymous):

\[x^2 + 10x + 25 - 25 = 15\] \[(x+5)^2 = (\sqrt{40})^2\]

OpenStudy (anonymous):

add 25 on both sides. it'll become (x+5)^2 = 40

Parth (parthkohli):

Here, \(2ab = 10x\) \( \color{Black}{\Rightarrow ab = 5x}\) \(b = 5\) \(a = x\)

OpenStudy (anonymous):

okay im bad at math so i need the right answer

OpenStudy (anonymous):

Yes... Or you can say that we add square of half the coefficient of x..

Parth (parthkohli):

Missing term is \(b^2\)\( \color{Black}{\Rightarrow (5)^2 }\)

OpenStudy (lgbasallote):

hah s many answers

OpenStudy (lgbasallote):

so*

OpenStudy (anonymous):

so its 25?

Parth (parthkohli):

Yep

OpenStudy (anonymous):

yes

Parth (parthkohli):

And you can even see that \(x^2 + 10x + 25 = (x + 5)^2\)

OpenStudy (anonymous):

Yes it is 25..

OpenStudy (anonymous):

we have to add 40

OpenStudy (anonymous):

thanks

Parth (parthkohli):

So it's a perfect square.

OpenStudy (anonymous):

Welcome dear..

Parth (parthkohli):

The final equation will then look like \(x^2 + 10x + 25 = 40\)

OpenStudy (anonymous):

what abpout What number must you add to complete the square? x^2 + 4x = 15

OpenStudy (anonymous):

Its 25!! You have to add 25!

OpenStudy (anonymous):

here a=2

Parth (parthkohli):

@LalanaG123 Can you please do it in a new post?

OpenStudy (anonymous):

Here what is the half of coefficient of x???